OpenStudy (anonymous):

if alpha and beta are the roots of 3x^2 - 4x+8 = 0 find the value of: alpha*beta

OpenStudy (anonymous):

For a quadratic equation, ax^2+bx+c \[\alpha \beta = c/a\]